x(dy)/(dx)=x^(2) + 5y

1
Expert's answer
2022-04-08T02:28:16-0400
"xy'=x^2+5y"

"y'-\frac{5}{x}y=x"

We put

"y=uv\\
y'=u'v+uv'"

Then

"u'v+uv'-\frac{5}{x}uv=x"

"u(v'-\frac{5}{x}v)+u'v=x"

Let

"v'-\frac{5}{x}v=0"

Then

"v'=\frac{5}{x}v"

"\ln v=5\ln x"

"v=x^5"

Thus

"x^5u'=x"

"u'=x^{-4}"

"u=-1\/3x^{-3}+C"

Finally

"y=uv=(-1\/3x^{-3}+C)x^5"

"y=-1\/3x^2+Cx^5"